### v1

@@ Abstract
Aims. We present a preview of the Stellar Stream Legacy Survey (SSLS) in the dwarf galaxy regime. The SSLS aims to produce a statistically large, homogeneous sample of stellar streams, in order to yield new constraints on galaxy formation through comparison with theory.
Methods. We visually inspect dwarf galaxies using the DESI Legacy Imaging Survey (DES and DECaLS footprints, -band29 mag arcsec-2) within 4 – 35 Mpc. We develop a classification metric to categorise accretion debris around dwarf galaxies, and measure the frequency of accretion features in the DES footprint.
Results. We present the first release of accretion features around dwarf galaxies, including 1 stream, 11 shells, and 8 asymmetric stellar halos, of which 17 constitute new identifications. We find that 5.07 of dwarfs (37/730) contain detectable accretion features, which is lower than the SSLS results for massive galaxies, and discuss a tentative shell detection bias.
Conclusions. Our results highlight the difficulty of detecting streams around dwarf galaxies, and identify the need for improved theoretical modelling of low-mass merger morphologies. Nevertheless, they place observational constraints on hierarchical mass assembly in this regime.
@@
While we know of at least 100+ extragalactic streams around MW–mass hosts, significantly less progress has been made in identifying streams around dwarf galaxies. Beyond our Local Group, only a handful of dwarfs with stellar streams are known (e.g., NGC 4449, DDO 68, NGC 300; Martínez-Delgado et al. 2012; Annibali et al. 2016; Fielder et al. 2025). Targetted campaigns, such as the Smallest Scale of Hierarchy survey (SSH; Annibali et al. 2020), are resolving accretion signatures in stellar halos up to the edge of the Local Volume (¡11 Mpc) in 45 isolated dwarfs (Annibali et al. 2022; Pascale et al. 2022, 2024; Sacchi et al. 2024). Yet the frequency of hierarchical mass assembly in the low-mass regime is not well understood. While previous works have catalogued merging dwarf galaxies (z ¡ 0.02, Paudel et al. 2018), or measured the frequency of tidal interaction debris around isolated dwarf galaxies (z ¡ 0.12, Kado-Fong et al. 2020), a frequency estimate following a careful identification of streams, shells and other accretion signatures (distinguished from tidal arms) has not yet been compiled, in order to examine CDM cosmological simulations with sufficient detail.
@@ 2 Data and method
Our aims consist of (i), a proof-of-concept pilot visual inspection of the SSLS footprint covered by the Dark Energy Camera (DES, DECaLS) to create a satellite accretion classification metric (objective 1), and (ii), obtaining the frequency of this event through a more advanced visual inspection of the DES footprint, guided by this metric (objective 2).
We curate two (DES, DECaLS) dwarf galaxy samples by crossmatching the 50 Mpc Galaxy Catalog (50MGC; Ohlson et al. 2024) to the footprints. The 50MGC catalogue provides homogenized mass, distance and morphological type measurements for 15,424 nearby galaxies from three sources: HyperLeda (Makarov et al. 2014), the Local Volume Galaxy (LVG) catalogue (Karachentsev et al. 2013) and the NASA-Sloan Atlas (NSA)111https://www.sdss.org/dr13/manga/manga-target-selection/nsa/. Dwarf galaxies are selected with stellar mass , has a distance between Mpc and galactic latitude , using the bestdist, logmass columns. For each dwarf, we calculate the projected separation and velocity separation (using the v_cmb column) from other galaxies in the 50MGC catalogue. Only dwarfs with kpc and km s-1 from a massive galaxy () are selected. Dwarf galaxy pairs are flagged ( kpc, km s-1) and discarded during the visual inspection (next step). This is not done automatically as we take care to not remove disrupted, smaller satellite progenitors that are part of an accretion feature. However, we do remove dwarfs with a smaller, intact satellite that has (at least) a distance measurement. We do not remove satellites with no catalogue entry, given it is unclear of their distance and association. Future SSLS releases will implement a tailored isolation criteria to fine-tune the frequency estimate for comparisons with cosmological simulations, once they become available. The sample is visually inspected for accretion features using the DESI Legacy Imaging Survey (DESI LS222https://www.legacysurvey.org/viewer) sky browser DR10 images. DESI LS includes other survey data that provide additional information on each galaxy (e.g., we used GALEX to check for recent star formation, sometimes indicative of tidal arms). We take special care to not mistake spiral arms as accretion features, e.g., bright blue arms extending from the galactic disc, or dusty spiral arms coalescing after dwarf-dwarf post-mergers.
@@ 3 Results and discussion
In Figure 2, we present the most spectacular accretion features around dwarf galaxies discovered in our proof-of-concept study (objective 1). We order them by apparent morphology and show a stream (ESO 508-059), asymmetric stellar halos (ESO 358-054 horizontally to UGC 6018) and shells (IC 700 through NGC 4928). 17 of these 20 cases are new discoveries, with only PGC 40604, ESO 149-003, and NGC 4765 previously known. We discuss each case individually in Appendix A.
@@
We find only one stream morphology in the entire survey (DECaLS footprint) and none in DES. Within the DES footprint, fewer than 5.07 (37) of dwarf galaxies show an accretion feature (shell/asymmetric stellar halo), including 15 very low-mass () cases. In the same DES footprint, and using the same data and SB limits, Miró-Carretero et al. (2024) found of MW-mass galaxies to have an accretion feature (stream, shell, and other morphologies). While this could indicate genuine differences for merger rates and debris morphologies across mass regimes, several observational factors complicate direct comparison. First, dwarf minor mergers (1:10 total mass) may not contribute enough stars to appear in stellar halo SB profiles. The stellar mass – halo mass relation at these scales is poorly known and depends on the assumed halo mass threshold for galaxy formation, which is not established. Major mergers (1:3 total mass) should be more detectable as, not only do they bring in more stars, but they can also form extended distributions in the stellar halo (due to ‘pulled-out’ host material) affecting SB profiles (Deason et al. 2022). Second, dwarf major mergers might preferentially form shells. While formation depends primarily on the progenitor orbit (radial versus circular/polar), there are secondary effects depending on mass. Amorisco (2017) has shown that dynamical friction can act to radialize orbits when the accreted dwarf is relatively massive. Scaling this to dwarf mergers, one could argue that major mergers are more likely to have radial orbits and create shells. Taken together, shells from major mergers could be more structurally stable, and thus detectable for longer, than streams. Indeed, the stability of streams formed by dwarf mergers, and how long they last before they phase mix and create asymmetric stellar halos (like those in Figure 2), is not yet constrained. Finally, the detectability and classification bias of streams versus shells should be explored, given we can only make classifications based on the brightest pieces of debris. Dwarf tidal debris is already at the observable limit (e.g., Deason et al. 2022; Tau et al. 2025), and the viewing angle is expected to affect the apparent SB and even classification— for instance, ‘edge-on’ streams will appear brighter (and potentially shell-like) in comparison to ‘face-on’ streams. The effect of the viewing angle on the detectability of streams will be explored as future work (Sakowska et al., in prep). In summary, this Letter highlights the difficulty in finding streams around dwarf galaxies, and finds tentative evidence for a shell detection bias.
@@
We found it challenging to unambiguously separate accretion from morphological features without further theoretical information on the dwarf-dwarf merger sequence. During the inspection we identified a variety of complex dwarf morphologies, including 20 dwarf post-mergers with spiral arms still mixing (which resembled accretion tracks). We discuss these observational challenges in Appendix B. Improving our statistic requires an extensive library of -body hydrodynamical simulations spanning the dwarf-dwarf merger sequence. Such simulations have successfully reproduced faint substructures seen in nearby (¡ 11 Mpc) isolated dwarf galaxies (e.g., DDO 68, NGC 5238, UGC 8760, Pascale et al. 2022, 2024) and 1:10 dwarf pairs (Besla et al. 2016; Pearson et al. 2018). Ideally, this library would also account for SB-limits and viewing angle effects on stream appearance and brightness.
@@ 4 Conclusions
This Letter comprises the first SSLS release of 20 dwarf galaxies with accretion features (1 stream, 11 shells, and 8 asymmetric stellar halos) identified by visual inspection of the DES/DECaLS footprints within 4 – 35 Mpc. Using the same data and SB-limits as for our previous studies on MW-like galaxies, we find that dwarf galaxies show less () accretion features than MW-like galaxies (). However, the observability of these extremely faint features must be investigated further before confirming a physical difference in the merger rates between these mass regimes. Specifically, we should assess whether we are observationally biased towards detecting primarily major mergers, and if this bias is an important driver behind the dominance of shell features in our survey. Additionally, an extensive -body hydrodynamical model library of dwarf-dwarf satellite mergers is needed to better disentangle accretion from morphological features.
Looking forward, our survey will provide dwarf galaxy stream candidates for follow-up studies- e.g. DM halo fitting, a technique expected to place novel constraints on DM candidates. The frequency estimate presented here will test predictions made by cosmological simulations, as they become more readily available in the low-mass regime. Finally, our work previews what will be achievable with upcoming wide-field surveys, i.e. Euclid, LSST and the Nancy Grace Roman Space Telescope, for discovering new stellar streams around dwarf galaxies.
@@ Acknowledgements.
We thank Francisco Prada and Justin Read for insightful discussions which helped improve the manuscript. JS acknowledges financial support from project PID2022-138896NB-C53 and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. DMD thanks financial support for a visiting researcher stay at the Astronomy and Astrophysics Department of the University of Valencia within the framework of the «Talent Attraction» programme implemented by the Office of the Vice-Principal for Research (INV25-01-15). TS gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. TS was supported by NSF through grant AST-2510183 and by NASA through grants 22-ROMAN22-0055 and 22-ROMAN22-0013.References
@@
Appendix A Challenges in identifying stellar streams around dwarf galaxies
Appendix B The diversity of satellite accretion in the dwarf galaxy regime

### v2

@@ Abstract
Aims. We present a preview of the Stellar Stream Legacy Survey (SSLS) in the dwarf galaxy regime. The SSLS aims to produce a statistically large, homogeneous sample of stellar streams for comparison with galaxy evolution theory.
Methods. We visually inspect dwarf galaxies using the DESI Legacy Imaging Survey (DES and DECaLS footprints, -band29 mag arcsec-2) within 4 – 35 Mpc. We develop a classification metric to categorise accretion debris around dwarf galaxies, and measure the frequency of accretion features in the DES footprint only.
Results. We present the first release of accretion features around dwarf galaxies collected from the DES and DECaLS footprints, including 1 stream, 11 shells, and 8 asymmetric stellar halos, of which 17 constitute new identifications. In the DES footprint, we inspect 730 dwarfs and find that 5.1 (37/730) show accretion features. Although this frequency measurement is lower than the SSLS result for massive galaxies, we discuss the observational biases behind detecting streams in the dwarf galaxy regime.
Conclusions. Our results highlight the difficulty of detecting streams around dwarfs, and identify the need for improved theoretical modelling of low-mass merger morphologies. Nevertheless, they place constraints on hierarchical mass assembly in this regime.
@@
While we know of at least 100+ extragalactic streams around MW–mass hosts, significantly less progress has been made in identifying streams around dwarf galaxies. Beyond our Local Group, only a handful of dwarfs with stellar streams are known (e.g., NGC 4449, DDO 68, NGC 300; Martínez-Delgado et al. 2012; Annibali et al. 2016; Fielder et al. 2025). Targetted campaigns, such as the Smallest Scale of Hierarchy survey (SSH; Annibali et al. 2020), are resolving accretion signatures in stellar halos up to the edge of the Local Volume (¡11 Mpc) in 45 isolated dwarfs (Annibali et al. 2022; Pascale et al. 2022, 2024; Sacchi et al. 2024). Yet the frequency of hierarchical mass assembly in the low-mass regime is not well understood. Previous works have catalogued merging dwarfs (z ¡ 0.02; Paudel et al. 2018), or measured the frequency of tidal debris around isolated dwarfs (z ¡ 0.12; Kado-Fong et al. 2020), but no frequency measurement of streams and shells (distinguished from tidal arms) has been presented. Such a result would enable cross-examination with future predictions made by cosmological simulations on detectable accretion features in the low-mass regime.
@@ 2 Data and method
Our aims consist of (i), a proof-of-concept pilot, manual visual inspection of the SSLS footprint covered by the Dark Energy Camera (DES, DECaLS) to create a satellite accretion classification metric (objective 1), and (ii), obtaining the frequency of this event through a quantitative, manual visual inspection of the DES footprint only, guided by this metric (objective 2).
We curate two (DES, DECaLS) dwarf galaxy samples by crossmatching the 50 Mpc Galaxy Catalog (50MGC; Ohlson et al. 2024) to the footprints. The 50MGC catalogue provides homogenized mass, distance and morphological type measurements for 15,424 nearby galaxies from three sources: HyperLeda (Makarov et al. 2014), the Local Volume Galaxy (LVG) catalogue (Karachentsev et al. 2013) and the NASA-Sloan Atlas (NSA)111https://www.sdss.org/dr13/manga/manga-target-selection/nsa/. Dwarf galaxies are selected with stellar mass , has a distance between Mpc and galactic latitude , using the bestdist, logmass columns. For each dwarf, we calculate the projected separation and velocity separation (using the v_cmb column) from other galaxies in the 50MGC catalogue. Only dwarfs with kpc and km s-1 from a massive galaxy () are selected. Dwarf galaxy pairs are flagged ( kpc, km s-1) and discarded during the visual inspection (next step). This is not done automatically as we take care to not remove disrupted, smaller satellite progenitors that are part of an accretion feature. However, we do remove dwarfs with a smaller, intact satellite that has (at least) a distance measurement. We do not remove satellites with no catalogue entry, given it is unclear of their distance and association. Future SSLS releases will implement a tailored isolation criteria to fine-tune the results for comparison with predictions of observable dwarf accretion features from cosmological simulations, akin to Miró-Carretero et al. (2025) for MW-mass hosts, once they become available. The sample is manually inspected for accretion features using the DESI Legacy Imaging Survey (DESI LS222https://www.legacysurvey.org/viewer) DR10 images. DESI LS includes other survey data that provide additional information on each galaxy (e.g., we used GALEX to check for recent star formation, sometimes indicative of tidal arms). We take care to not mistake spiral arms as accretion features, e.g., bright blue arms extending from the galactic disc, or dusty spiral arms coalescing after dwarf post-mergers.
@@ 3 Results and discussion
In Figure 2, we select 10 positive cases each from DES and DECaLS to present the most spectacular examples of accretion features discovered in our proof-of-concept study (objective 1). We order them by apparent morphology and show a stream (ESO 508-059), asymmetric stellar halos (ESO 358-054 horizontally to UGC 6018) and shells (IC 700 through NGC 4928). 17 of these 20 cases are new discoveries, with only PGC 40604, ESO 149-003, and NGC 4765 previously known. We discuss each case individually in Appendix B.
@@
We find only one stream morphology in the entire survey (DECaLS footprint) and none in DES. Within the DES footprint, 5.1 (37) of dwarf galaxies show an accretion feature (shell/asymmetric stellar halo), including the 10 shown in Figure 2. We find 15 very low-mass () cases. In the same DES footprint, and using the same data and SB limits, Miró-Carretero et al. (2024) found of MW-mass galaxies to have an accretion feature (stream, shell, and other morphologies). While this could indicate genuine differences for merger rates and debris morphologies across mass regimes, several observational factors complicate direct comparison. First, dwarf minor mergers (1:10 total mass) may not contribute enough stars to appear in stellar halo SB profiles. The stellar mass – halo mass relation at these scales is poorly known and depends on the assumed halo mass threshold for galaxy formation, which is not established. Major mergers (1:3 total mass) should be more detectable as, not only do they bring in more stars, but they can also form extended distributions in the stellar halo (due to ‘pulled-out’ host material) affecting SB profiles (Deason et al. 2022). Second, dwarf major mergers might preferentially form shells. While formation depends primarily on the progenitor orbit (radial versus circular/polar), there are secondary effects depending on mass. Amorisco (2017) has shown that dynamical friction can act to radialize orbits when the accreted dwarf is relatively massive. Scaling this to dwarf mergers, one could argue that major mergers are more likely to have radial orbits and create shells. Taken together, shells from major mergers could be more structurally stable, and thus detectable for longer, than streams. Indeed, the stability of streams formed by dwarf mergers, and how long they last before they phase mix and create asymmetric stellar halos (like those in Figure 2), is not yet constrained. Finally, the detectability and classification bias of streams versus shells should be explored, given we can only make classifications based on the brightest pieces of debris. Dwarf tidal debris is already at the observable limit (e.g., Deason et al. 2022; Tau et al. 2025), and the viewing angle is expected to affect the apparent SB and even classification- for instance, ‘edge-on’ streams will appear brighter (and potentially shell-like) in comparison to ‘face-on’ streams. The effect of the viewing angle on the detectability of streams will be explored as future work (Sakowska et al., in prep). In summary, we highlight the difficulty in finding streams around dwarfs, and find tentative evidence for a shell detection bias.
@@
We found it challenging to unambiguously separate accretion from morphological features without further theoretical information on the dwarf-dwarf merger sequence. During the inspection we identified a variety of complex dwarf morphologies, including 20 dwarf post-mergers with spiral arms still mixing (which resembled accretion tracks). We discuss these observational challenges in Appendix A. For this reason, we consider 5.1 to be an upper limit to the observable number of accretion features in the DES sample at this SB-limit. Improving our statistic requires an extensive library of -body hydrodynamical simulations spanning the dwarf-dwarf merger sequence. Such simulations have successfully reproduced faint substructures seen in nearby (¡ 11 Mpc) isolated dwarf galaxies (e.g., DDO 68, NGC 5238, UGC 8760, Pascale et al. 2022, 2024) and 1:10 dwarf pairs (Besla et al. 2016; Pearson et al. 2018). Ideally, this library would also account for SB-limits and viewing angle effects on stream appearance and brightness.
@@ 4 Conclusions
This Letter comprises the first SSLS release of 20 dwarf galaxies with accretion features (1 stream, 11 shells, and 8 asymmetric stellar halos) identified by visual inspection of the DES/DECaLS footprints within 4 – 35 Mpc. In the DES footprint only, using the same data and SB-limits as for our previous studies, we find that dwarfs show comparatively less () accretion features than MW-like galaxies (). An extensive -body hydrodynamical model library of dwarf-dwarf satellite mergers is needed to better disentangle accretion from morphological features and constrain our result, which we regard as an upper limit. The observability of these extremely faint features must be investigated further before confirming a physical difference in the merger rates. We should assess whether we are observationally biased towards detecting major mergers, and if this is an important driver behind the dominance of shells in our survey. Looking forward, our survey will provide dwarf galaxy stream candidates for follow-up studies- e.g. DM halo fitting, a technique expected to place novel constraints on DM candidates. The frequency measurement provided will test upcoming predictions made by cosmological simulations on observable accretion features in low-mass galaxies. Finally, our work previews what will be achievable with upcoming wide-field surveys, i.e. Euclid, LSST and the Nancy Grace Roman Space Telescope, for discovering new stellar streams around dwarf galaxies.
@@ Acknowledgements.
We thank the anonymous referee, F. Prada and J. Read for insightful comments which helped improve the manuscript. JS acknowledges financial support from project PID2022-138896NB-C53 and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. DMD thanks financial support for a visiting researcher stay at the Astronomy and Astrophysics Department of the University of Valencia within the framework of the «Talent Attraction» programme implemented by the Office of the Vice-Principal for Research (INV25-01-15). TS gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. TS was supported by NSF through grant AST-2510183 and by NASA through grants 22-ROMAN22-0055 and 22-ROMAN22-0013.References
@@
Appendix A The diversity of satellite accretion in the dwarf galaxy regime
@@
Appendix B Challenges in identifying stellar streams around dwarf galaxies